# Proactive Recommendation in Social Networks: Steering User Interest via Neighbor Influence

### v1

@@ Abstract.
Recommending items solely catering to users’ historical interests narrows users’ horizons. Recent works have considered steering target users beyond their historical interests by directly adjusting items exposed to them. However, the recommended items for direct steering might not align perfectly with users’ interests evolution, detrimentally affecting target users’ experience.
@@
To avoid this issue, we propose a new task named Proactive Recommendation in Social Networks (PRSN) that indirectly steers users’ interest by utilizing the influence of social neighbors, i.e., indirect steering by adjusting the exposure of a target item to target users’ neighbors. The key to PRSN lies in answering an interventional question: what would a target user’s feedback be on a target item if the item is exposed to the user’s different neighbors? To answer this question, we resort to causal inference and formalize PRSN as: (1) estimating the potential feedback of a user on an item, under the network interference by the item’s exposure to the user’s neighbors; and (2) adjusting the exposure of a target item to target users’ neighbors to trade off steering performance and the damage to the neighbors’ experience. To this end, we propose a Neighbor Interference Recommendation (NIRec) framework with two key modules: (1) an interference representation-based estimation module for modeling potential feedback; and (2) a post-learning-based optimization module for optimizing a target item’s exposure to trade off steering performance and the neighbors’ experience by greedy search. We conduct extensive semi-simulation experiments based on three real-world datasets, validating the steering effectiveness of NIRec.
Proactive Recommendation, Causal Recommendation, Causal Inference under Network Interference
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. INTRODUCTION
@@
Figure 1. Steering Recommendation. (a) The target user does not watch Shooting Fish exposed to him/her directly. (b) The target user watches Little City but the interests do not evolve with the item. (c) The target item is exposed to the target user’s neighbors with the corresponding interests, leading to a successful steer.
@@
Recommender systems are widely used in various fields such as social media (Zhang et al., 2023c), content sharing (Gao et al., 2023; Zhang et al., 2023a), and e-commerce (Wei et al., 2023). A de facto paradigm of recommender models is to learn users’ interests from their historical feedback for personalized recommendation (Gao et al., 2022). However, reinforcing historical interests leads to filter bubbles, restricting users’ exposure to diverse viewpoints and exacerbating issues like opinion polarization (Gao et al., 2024; Wang et al., 2022). Therefore, it is crucial to build proactive recommendation that steer users beyond their historical interests.
@@
Existing studies on proactive recommendation directly adjust the exposed items to target users to steer user interests, roughly falling into two research lines: steering without a target item (Carbonell and Goldstein, 1998; Chen et al., 2018; Wang et al., 2022; Yang et al., 2023b) and steering with a target item (Zhu et al., 2023; Bi et al., 2024). The former line directly exposes diverse items to target users to broaden users’ horizons beyond historical interests (Carbonell and Goldstein, 1998; Chen et al., 2018; Wang et al., 2022; Yang et al., 2023b). However, they overlook users’ willingness to interact with those items, potentially leading to failure in interest steering (see Figure 1 (a)). The latter line proactively steers a target user to interact with a target item by progressively exposing a sequence of carefully chosen items to the target user, gradually bridging the target user’s current interests with the target item (Zhu et al., 2023; Bi et al., 2024). However, crafting such a sequence is significantly challenging, since it might not align perfectly with the user’s interest evolution (see Figure 1 (b)). Misjudging the trajectory of the user’s interest evolution could lead to inappropriate recommendations and detrimentally affect the user’s experience.
@@
Figure 2. Causal Graphs. (a) The outcome of the unit is only influenced by its features and treatment. (b) The unit’s outcome is influenced by the neighbors’ features and treatments.
@@
To avoid the problems of directly steering, we introduce a new task named Proactive Recommendation in Social Networks (PRSN) — adjusting a target item’s exposure to target users’ neighbors to steer their interests indirectly, given users’ feedback on a target item is influenced by their neighbors on a social network (Anagnostopoulos et al., 2008; Lewis et al., 2012). An illustration is shown in Figure 1 (c), a target user without interest in romance movies is nudged to watch Touch by the user’s neighbors who enjoy this romance movie. PRSN adjusts a target item’s exposure to the neighbors with two objectives: (1) improving interaction probability between target users and the target item, and (2) limiting the damage to the neighbors’ experience. The main consideration for the second objective is the potential compromise of neighbors’ experience when increasing the target item’s exposure. For instance, to pursue steering, the target item might displace another item that better aligns with the neighbors’ interests.
@@
To achieve the objectives, the key lies in answering an interventional question: what would a target user’s feedback be on a target item if the item is exposed to the user’s different neighbors? To answer this question, we resort to causal inference to estimate the potential outcomes under network interference (Ogburn and VanderWeele, 2014; Ma and Tresp, 2021; Ma et al., 2022). However, the traditional potential outcome framework for recommendation typically assumes no interference across units, where each unit is a user-item pair $(u,i)$. As shown in Figure 2(a), the outcome $Y_{ui}$ of whether user $u$ interacts with item $i$ is only influenced by the treatment of whether exposing item $i$ to user $u$ and the features of the unit. To model the influence of the neighbors on target users, we relax the no-interference assumption in Figure 2(b) and estimate target users’ potential feedback on a target item interfered by the target item’s exposure to users’ neighbors. The new causal paths from the neighbor units’ treatments and features to $Y_{ui}$ indicate that the effects of neighbor units spill over to the target unit in social networks, interfering with the outcome of the target unit. This interference-aware potential framework shows that the essence of achieving PRSN lies in: (1) estimating interference-aware potential outcomes for each unit; and (2) assigning treatment values of the neighbor user-item units to improve the probabilities of positive potential outcome of target user-item unit while preventing the damage to the neighbors’ experience, as guided by the estimated potential outcomes.
@@
To this end, we propose a Neighbor Interference Recommendation (NIRec) framework, which consists of two key modules: (1) an interference representation-based estimation module; and (2) a post-learning-based optimization module. The interference representation-based estimation module integrates a unit’s individual representation with its neighbor units’ interference representations to obtain the interference-aware potential outcome. The unit’s individual representation captures the user’s individual interests regarding the item. Meanwhile, the interference representations measure the interference effects of neighbor units on the target unit. NIRec obtains interference representations by aggregating the neighbor units’ treatments and individual representations by masked graph convolution with attention. Given the well-trained estimation module and a pre-designed user experience cost function, the post-learning-based module greedily searches the values of neighbors’ treatments to jointly maximize the positive potential outcomes probabilities of target units and minimize the decrease of the expected ratings of the neighbors. We conduct extensive semi-simulation experiments on real-world datasets with in-depth analysis under various settings (e.g., various target users and target items), validating the steering effectiveness of NIRec.
@@
  • We first propose a novel PRSN task, which adjusts a target item’s exposure to target users’ neighbors to indirectly steer target users’ interests. This task emphasizes the significance of leveraging neighbors in social networks to expand users’ interests.
  • We design NIRec to achieve PRSN, which leverages an interference representation-based estimation module to estimate interference-aware potential outcomes and a post-learning-based optimization module to optimize treatment values to trade off steering performance and the experience of target users’ neighbors.
  • We conduct substantial semi-simulation experiments on real-world datasets, verifying the effectiveness of NIRec.
2. TASK FORMALIZATION
@@
2.1. A Causal View of Traditional
Recommendation
@@ 2.1. A Causal View of Traditional
In the factual world, we can only observe one potential outcome of unit $(u,i)$ under one treatment value, which is called the outcome of unit $(u,i)$ and denoted as $Y_{ui}$. For example, if unit $(u,i)$ was assigned with a treatment value of $O_{ui}=0$, potential outcome $Y_{ui}(O_{ui}=1)$ cannot be observed. Correspondingly, the observed feedback data in the factual world is denoted as $\mathcal{D}=\{\bm{x}_{ui},o_{ui},y_{ui}|o_{ui}=1,u\in\mathcal{U},i\in\mathcal{I}\}$. The objective of traditional recommendation is to: (1) estimate the probability of positive potential outcome $P(Y_{ui}(O_{ui}=1)=1|\bm{X}_{ui}=\bm{x}_{ui})$ for each unit by the observed feedback data; and (2) recommend items with the highest probabilities to cater users.
2.2. Formalization of PRSN with Causal Inference Under Network Interference
@@
Considering that users’ feedback on a target item is influenced by their neighbors in social networks (Anagnostopoulos et al., 2008; Lewis et al., 2012), we believe that the interests of a user’s neighbors in the exposed items can be propagated through social networks. Therefore, we propose PRSN, which indirectly nudges target users to interact with a target item by adjusting this item’s exposure to the social neighbors of target users. Given that increasing the target item’s exposure may damage the neighbors’ experience, PRSN has two objectives: (1) improving the interaction probability between the target users and the target item; and (2) limiting the damage to the neighbors’ experience. The key to PRSN is to answer an interventional question: what would a target user’s feedback on a target item be if the item is exposed to the user’s different neighbors? To answer this question, we resort to causal inference under network interference (Ogburn and VanderWeele, 2014; Ma and Tresp, 2021; Ma et al., 2022).
@@
Let $\mathcal{S}=\{(u,u^{\prime})\mid u\in\mathcal{U},u^{\prime}\in\mathcal{U},s_{uu^{\prime}}=1\}$ denote a directed social network, where $s_{uu^{\prime}}=1$ indicates that user $u$ trusts user $u^{\prime}$. We define user $u$’s neighbors as the users connected with user $u$ in social network $\mathcal{S}$, denoted as $\mathcal{N}_{u}=\{u^{\prime}|s_{uu^{\prime}}=1,u^{\prime}\in\mathcal{U}\}$. Similarly, we define the neighbor units of unit $(u,i)$ as the user-item pairs composed of the neighbors of user $u$ and item $i$, which is denoted as $(\mathcal{N}_{u},i)=\{(u^{\prime},i)\mid u^{\prime}\in\mathcal{N}_{u}\}$. Treatments of unit $(u,i)$’s neighbor units are denoted as $\bm{O}_{\mathcal{N}_{u}i}=[O_{u^{\prime}_{1}i},O_{u^{\prime}_{2}i},\cdots,O_{u^{\prime}_{|\mathcal{N}_{u}|}i}]^{\intercal}$. Furthermore, $\mathcal{U}^{\ast}\in\mathcal{U}$ and $i^{\ast}\in\mathcal{I}$ are target users and a target item, respectively. $\mathcal{U}^{\ast}$’s neighbors are users connected to them in the social network while excluding those who belong to $\mathcal{U}^{\ast}$, which is denoted as $\mathcal{N}_{\mathcal{U}^{\ast}}=\{u^{\prime}|s_{uu^{\prime}}=1,u\in\mathcal{U}^{\ast},u^{\prime}\notin\mathcal{U}^{\ast}\}$. With the presence of neighbor influence in social networks, it is necessary to be aware of two types of new causal paths related to network interference (see Figure 2(b)). In particular, $\bm{X}_{u^{\prime}i},O_{u^{\prime}i}\rightarrow Y_{ui}$ depicts that the effect of neighbor $(u^{\prime},i)$’s treatment $O_{u^{\prime}i}$ and features $\bm{X}_{u^{\prime}i}$ spills over to unit $(u,i)$, interfering with unit $(u,i)$’s outcome $Y_{ui}$. It relaxes the no-interference assumption of traditional causal recommendation.
@@
  • Estimating the probability of a positive potential outcome under varying interference (i.e., $\bm{o}_{\mathcal{N}_{u}i}$) for each unit:
    (1) $P\left(Y_{ui}\left(O_{ui}=1,\bm{O}_{\mathcal{N}_{u}i}=\bm{o}_{\mathcal{N}_{u}i}\right)=1\mid\bm{X}_{ui}\right),$
@@
  • Optimizing $\bm{o}_{\mathcal{N}_{\mathcal{U}^{\ast}}i^{\ast}}$ to trade off the sum of probabilities of positive potential outcomes of target units $\sum_{u^{\ast}\in\mathcal{U^{\ast}}}P(Y_{u^{\ast}i^{\ast}}(\bm{O}_{\mathcal{U}^{\ast}i^{\ast}}=\bm{1},\bm{O}_{\mathcal{N}_{\mathcal{U}^{\ast}}i^{\ast}}=\bm{o}_{\mathcal{N}_{\mathcal{U}^{\ast}}i^{\ast}})=1\mid\bm{X}_{u^{\ast}i^{\ast}}=\bm{x}_{u^{\ast}i^{\ast}})$ and the damage to the neighbors $\sum_{u^{\prime}\in\mathcal{N}_{\mathcal{U}^{\ast}}}c(o_{u^{\prime}i^{\ast}})$. Here $c(o_{u^{\prime}i^{\ast}})$ is a cost function measuring the damage of $o_{u^{\prime}i^{\ast}}$ to neighbor $u^{\prime}$ through the decrease in $u^{\prime}$ expected rating, with further details provided in Section 3.3.
3. METHOD
In this section, we introduce NIRec as shown in Figure 3, which consists of an interference representation-based potential outcome estimation module and a post-learning-based treatment values optimization module to achieve PRSN.
Figure 3. NIRec. Initially, train an interference representation-based estimation module for potential outcome estimation. Then, optimize the exposure of the target item to the user’s neighbor, given the well-trained estimation module and a pre-designed cost function.
3.1. Identification
We identify interference-aware potential outcomes with standard assumptions in the causal inference literature (Ogburn and VanderWeele, 2014; Ma and Tresp, 2021; Ma et al., 2022).
Assumption 1.
There exists an interference representation function $\phi(\cdot)$ sufficiently capturing information of $(u,i)$’s neighbor’ treatment values $\bm{o}_{\mathcal{N}_{u}i}$ and features $\bm{\mathrm{X}}_{\mathcal{N}_{u}i}$. If $\phi(\cdot)$ is determined, $(u,i)$’s potential outcome is determined. Formally, for any $(u,i)$, $\bm{o}^{1}_{\mathcal{N}_{u}i}$, $\bm{o}^{2}_{\mathcal{N}_{u}i}$, $\bm{\mathrm{X}}^{1}_{\mathcal{N}_{u}i}$, and $\bm{\mathrm{X}}^{2}_{\mathcal{N}_{u}i}$, if $\phi(\bm{o}^{1}_{\mathcal{N}_{u}i},\bm{\mathrm{X}}^{1}_{\mathcal{N}_{u}i})=\phi(\bm{o}^{2}_{\mathcal{N}_{u}i},\bm{\mathrm{X}}^{2}_{\mathcal{N}_{u}i})$, we have:
(2) $Y_{ui}\left(O_{ui}=1,\bm{O}_{\mathcal{N}_{u}i}=\bm{o}^{1}_{\mathcal{N}_{u}i}\right)=Y_{ui}\left(O_{ui}=1,\bm{O}_{\mathcal{N}_{u}i}=\bm{o}^{2}_{\mathcal{N}_{u}i}\right).$

### v2

@@ Abstract
Recommending items that solely cater to users’ historical interests narrows users’ horizons. Recent works have considered steering target users beyond their historical interests by directly adjusting items exposed to them. However, the recommended items for direct steering might not align perfectly with the evolution of users’ interests, detrimentally affecting the target users’ experience.
@@
To avoid this issue, we propose a new task named Proactive Recommendation in Social Networks (PRSN) that indirectly steers users’ interest by utilizing the influence of social neighbors, i.e., indirect steering by adjusting the exposure of a target item to target users’ neighbors. The key to PRSN lies in answering an interventional question: what would a target user’s feedback be on a target item if the item is exposed to the user’s different neighbors? To answer this question, we resort to causal inference and formalize PRSN as: (1) estimating the potential feedback of a user on an item, under the network interference by the item’s exposure to the user’s neighbors; and (2) adjusting the exposure of a target item to target users’ neighbors to trade off steering performance and the damage to the neighbors’ experience. To this end, we propose a Neighbor Interference Recommendation (NIRec) framework with two modules: (1) an interference representation-based estimation module for modeling potential feedback; (2) a post-learning-based optimization module for adjusting a target item’s exposure to trade off steering performance and the neighbors’ experience through greedy search. We conduct extensive semi-simulation experiments on real-world datasets, validating the steering effectiveness of NIRec. The code is available at https://github.com/HungPaan/NIRec.
keywords:
[…]
1 INTRODUCTION
@@
Figure 1: Steering Recommendation. (a) The target user does not watch Shooting Fish exposed to him/her directly. (b) The target user watches Little City but the interests do not evolve with the item. (c) The target item is exposed to the target user’s neighbors with the corresponding interests, leading to a successful steer.
@@
Recommender systems are widely used in various fields such as social media [zhang2023social], content sharing [gao2023rlrec], and e-commerce [wei2023longtail]. A de facto paradigm of recommender models is to learn users’ interests from their historical feedback for personalized recommendation [He2020lightgcn]. However, reinforcing historical interests leads to filter bubbles, restricting users’ exposure to diverse viewpoints and exacerbating issues like opinion polarization [gao2024filterbubble]. Therefore, it is crucial to build proactive recommendation that steers users beyond their historical interests.
@@
Existing studies on proactive recommendation directly adjust the exposed items to target users to steer user interests, roughly falling into two research lines: steering without a target item [car1998mrr, chen2018dpp, yang2023diversity] and steering with a target item [zhu2023influ, bi2024proactive]. The former line directly exposes diverse items to target users to broaden users’ horizons beyond historical interests [car1998mrr, chen2018dpp, yang2023diversity]. However, they overlook users’ willingness to interact with those items, potentially leading to failures in interest steering (see Figure 1 (a)). The latter line proactively steers a target user to interact with a target item by progressively exposing a sequence of carefully chosen items to the target user, gradually bridging the target user’s current interests with the target item [zhu2023influ, bi2024proactive, lian2025itmprec, wang2024incorporate]. However, crafting such a sequence is significantly challenging, since it might not align perfectly with the user’s interest evolution (see Figure 1 (b)). Misjudging the trajectory of the user’s interest evolution could lead to inappropriate recommendations and detrimentally affect the user’s experience.
@@
Figure 2: Causal Graphs. (a) The outcome of the unit is only influenced by its features and treatment. (b) The unit’s outcome is influenced by the neighbors’ features and treatments.
@@
To avoid the problems of directly steering, we introduce a new task named Proactive Recommendation in Social Networks (PRSN) — adjusting a target item’s exposure to target users’ neighbors to steer their interests indirectly, given users’ feedback on a target item is influenced by their neighbors on a social network [Aris2008social, lewis2012social]. An illustration is shown in Figure 1 (c), where a target user without interest in romance movies is nudged to watch Touch by the user’s neighbors who enjoy this romance movie. PRSN adjusts a target item’s exposure to the neighbors with two objectives: (1) improving interaction probability between target users and the target item, and (2) limiting the damage to the neighbors’ experience. The main consideration for the second objective is the potential compromise of neighbors’ experience when increasing the target item’s exposure. For instance, to pursue steering, the target item might displace another item that better aligns with the neighbors’ interests.
@@
To achieve the objectives, the key lies in answering an interventional question: what would a target user’s feedback be on a target item if the item is exposed to the user’s different neighbors? To answer this question, we need to resort to causal inference to estimate the potential outcomes under network interference [Ogburn2014caudiag, ma2021interference, ma2022hypergraph]. However, traditional causal recommendation frameworks [wang2019doub, ha2024finegrained] shown in Figure 2(a) typically assume no interference between units (i.e., user-item pairs). Specifically, outcome $Y_{ui}$, representing whether user $u$ interacts with item $i$, is influenced only by the treatment of item exposure to $u$ and the unit’s features, which can not answer the interventional question. To address this, we relax the no-interference assumption in Figure 2(b), where new causal paths from neighbor units’ treatments and features to $Y_{ui}$ indicate that the effects of neighbor units spill over, interfering with the target unit’s outcome. This interference-aware potential framework shows that the essence of achieving PRSN lies in: (1) estimating interference-aware potential outcomes for each unit; and (2) adjusting treatment values to neighbor units to trade off the probability of a positive potential outcome for the target unit and the damage to neighbors’ experience, as guided by the estimated potential outcomes.
@@
To this end, we propose a Neighbor Interference Recommend-ation (NIRec) framework, which consists of two key modules: (1) an interference representation-based estimation module; and (2) a post-learning-based optimization module. The interference representation-based estimation module integrates a unit’s individual representation with its neighbor units’ interference representations to obtain the interference-aware potential outcome. The unit’s individual representation captures the user’s individual interests regarding the item. Meanwhile, the interference representations measure the interference effects of neighbor units on the target unit. NIRec obtains interference representations by aggregating the neighbor units’ treatments and individual representations by masked graph convolution with attention. Given the well-trained estimation module and a pre-designed user experience cost function, the post-learning-based module greedily searches the values of neighbors’ treatments to jointly maximize the positive potential outcomes probabilities of target units and minimize the decrease of the expected ratings of the neighbors. We conduct extensive semi-simulation experiments on real-world datasets with in-depth analysis under various settings (e.g., various target users and target items), validating the steering effectiveness of NIRec.
@@
  • We propose a novel Proactive Recommendation in Social Networks (PRSN) task — adjusting a target item’s exposure to target users’ neighbors to indirectly steer target users’ interests. This task is orthogonal to existing research on direct steering, offering a new direction for steering user interest while avoiding the issues of direct steering.
  • We design a Neighbor Interference Recommendation (NIRec) framework to achieve PRSN, which leverages an interference representation-based estimation module to estimate interference-aware potential outcomes and a post-learning-based optimization module to adjust treatment values to trade off steering performance and the experience of target users’ neighbors.
  • We conduct semi-simulation experiments on real-world datasets with in-depth analysis under various settings, verifying the effectiveness of NIRec.
2 RELATED WORK
2.1 Steering Recommendation
The first line steers users without a target item, which directly improves recommendation diversity [car1998mrr, chen2018dpp, zheng2021diversity, yang2023diversity, zhang2023diversity]. For example,  [car1998mrr] reduces the redundancy of the recommendation list using greedy search-based re-ranking.  [zhang2023diversity] recommends multiple categories of items to users through disentangled representation learning. While these methods broaden users’ exposure to items beyond their historical interests, they do not account for users’ willingness to engage with those items, which limits their ability to effectively steer user interests. The other line steers with a target item [zhu2023influ, bi2024proactive]. They sequentially introduce intermediate items to users, bridging their existing interests with target items to steer their interest [zhu2023influ, bi2024proactive, wang2024incorporate, lian2025itmprec]. IRS [zhu2023influ] uses transformer-based models to create influence paths that lead users to predefined target items. IPG [bi2024proactive] enhances adaptability by dynamically adjusting strategies based on real-time user feedback. LLM-IPP [wang2024incorporate] explores integrating large language models’ reasoning and planning into proactive recommendation. Additionally, ITMPRec [lian2025itmprec] employs a multi-round strategy to proactively steer user interests toward target items. NIRec adjusts a target item’s exposure to a user’s neighbors to indirectly steer their interests, which is orthogonal to direct proactive recommendation, avoiding their issues while providing a new direction for steering user interests.
2.2 Causal Recommendation
One widely discussed task is to use a potential outcome framework to address the problem of missing not at random in recommendation [wang2019doub, guo2021mrdr, yang2023balance, zhou2023generalized, ha2024finegrained]. These methods rely on the no-interference assumption and focus on the unbiased learning of potential outcomes for all user-item pairs with exposure treatment values of 1.  [schnabel2016recommendations, imbens2015causal, zhou2023generalized] use various strategies to estimate inverse propensity scores and reweight the biased data to align the distribution of unbiased data.  [jiang2016doubly, wang2019doub, chen2021autodebias, guo2021mrdr, zhou2023generalized, ha2024finegrained] incorporate loss imputation with inverse propensity scores to achieve double robustness for unbiased recommendation. Causal recommendation also involves other problems with no-interference assumption [wu2022caurec]. For example,  [wu2022caurec] formulates the effect of position on user feedback within a potential outcome framework by considering the item’s position in the recommendation list as a treatment.  [sato2019uplift, sato2020uplift, xie2021causalcf] consider the scenario where a user can interact with an item without recommendation and estimate the uplift of the user’s feedback from recommendation by the potential outcome framework. Unlike them, our work relaxes the no-interference assumption and focuses on modeling interference-aware potential outcomes.
2.3 Causal Inference under Network Interference
Network interference receives widespread attention in causal inference for estimating individual treatment effects [Ogburn2014caudiag, ma2021interference, forastiere2021identification, ma2022hypergraph, jiang2022estimating, zhao2022interference]. For example,  [Ogburn2014caudiag] distinguishes between different causal mechanisms that give rise to interference using causal graphs.  [ma2022hypergraph] learns interference representation by hypergraph convolution with attention. Considering interference,  [ma2021interference] maximizes the utility on the entire network by optimizing all units’ treatments in the population on the network. Our work differs from these studies in two key ways: (1) constructing target user-item pairs and adjusting the neighbors’ treatment values to achieve our task for those target pairs; and (2) considering user experience under treatment adjustment.
3 TASK FORMALIZATION
@@
3.1 A Causal View of Traditional Recommendation
@@
In the factual world, we can only observe one potential outcome of unit $(u,i)$, which is called the outcome of unit $(u,i)$ and denoted as $Y_{ui}$. For example, if unit $(u,i)$ was assigned with a treatment value of $O_{ui}=0$, potential outcome $Y_{ui}(O_{ui}=1)$ cannot be observed. Correspondingly, the observed feedback data in the factual world is denoted as $\mathcal{D}=\{\bm{x}_{ui},o_{ui},y_{ui}|o_{ui}=1,u\in\mathcal{U},i\in\mathcal{I}\}$. The objective of traditional recommendation is to: (1) estimate the probability of positive potential outcome $P(Y_{ui}(O_{ui}=1)=1|\bm{X}_{ui}=\bm{x}_{ui})$ for each unit by the observed feedback data; and (2) recommend items with the highest probabilities to cater to users.
3.2 Formalization of PRSN with Causal Inference
Under Network Interference
@@ 3.2 Formalization of PRSN with Causal Inference
Considering that users’ feedback on a target item is influenced by their neighbors in social networks [Aris2008social, lewis2012social], we believe that the interests of a user’s neighbors in the exposed items can be propagated through social networks. Therefore, we propose PRSN, which indirectly nudges target users to interact with a target item by adjusting this item’s exposure to the social neighbors of target users. Given that increasing the target item’s exposure may damage the neighbors’ experience, PRSN has two objectives: (1) improving the interaction probability between the target users and the target item; and (2) limiting the damage to the neighbors’ experience. The key to PRSN is to answer an interventional question: what would a target user’s feedback on a target item be if the item is exposed to the user’s different neighbors? To answer this question, we resort to causal inference under network interference [Ogburn2014caudiag, ma2021interference, ma2022hypergraph].
@@
Let $\mathcal{S}=\{(u,u^{\prime})\mid u\in\mathcal{U},u^{\prime}\in\mathcal{U},s_{uu^{\prime}}=1\}$ denote a directed social network, where $s_{uu^{\prime}}=1$ indicates that user $u$ trusts user $u^{\prime}$. We define user $u$’s neighbors as the users connected with user $u$ in social network $\mathcal{S}$, denoted as $\mathcal{N}_{u}=\{u^{\prime}|s_{uu^{\prime}}=1,u^{\prime}\in\mathcal{U}\}$. Similarly, we define the neighbor units of unit $(u,i)$ as the user-item pairs composed of the neighbors of user $u$ and item $i$, which is denoted as $(\mathcal{N}_{u},i)=\{(u^{\prime},i)\mid u^{\prime}\in\mathcal{N}_{u}\}$. Treatments of unit $(u,i)$’s neighbor units are denoted as $\bm{O}_{\mathcal{N}_{u}i}=[O_{u^{\prime}_{1}i},O_{u^{\prime}_{2}i},\cdots,O_{u^{\prime}_{|\mathcal{N}_{u}|}i}]^{\intercal}$. Furthermore, $\mathcal{U}^{\ast}\subset\mathcal{U}$ and $i^{\ast}\in\mathcal{I}$ are target users and a target item, respectively. $\mathcal{U}^{\ast}$’s neighbors are users connected to them in the social network while excluding those who belong to $\mathcal{U}^{\ast}$, which is denoted as $\mathcal{N}_{\mathcal{U}^{\ast}}=\{u^{\prime}|s_{uu^{\prime}}=1,u\in\mathcal{U}^{\ast},u^{\prime}\notin\mathcal{U}^{\ast}\}$. With the presence of neighbor influence in social networks, it is necessary to be aware of two types of new causal paths related to network interference (see Figure 2(b)). In particular, $\bm{X}_{u^{\prime}i},O_{u^{\prime}i}\rightarrow Y_{ui}$ depicts that the effect of neighbor $(u^{\prime},i)$’s treatment $O_{u^{\prime}i}$ and features $\bm{X}_{u^{\prime}i}$ spills over to unit $(u,i)$, interfering with unit $(u,i)$’s outcome $Y_{ui}$. It relaxes the no-interference assumption of traditional causal recommendation.
@@
  • Estimating the probability of positive potential outcomes under varying interference (i.e., $\bm{o}_{\mathcal{N}_{u}i}$) for each unit:
    $P\left(Y_{ui}\left(O_{ui}=1,\bm{O}_{\mathcal{N}_{u}i}=\bm{o}_{\mathcal{N}_{u}i}\right)=1\mid\bm{X}_{ui}\right),$ (1)
@@
  • Balancing the sum of probabilities of positive potential outcomes of target units $\sum_{u^{\ast}\in\mathcal{U^{\ast}}}P(Y_{u^{\ast}i^{\ast}}(\bm{O}_{\mathcal{U}^{\ast}i^{\ast}}=\bm{1},\bm{O}_{\mathcal{N}_{\mathcal{U}^{\ast}}i^{\ast}}=\bm{o}_{\mathcal{N}_{\mathcal{U}^{\ast}}i^{\ast}})=1\mid\bm{X}_{u^{\ast}i^{\ast}}=\bm{x}_{u^{\ast}i^{\ast}})$ and the damage to the neighbors $\sum_{u^{\prime}\in\mathcal{N}_{\mathcal{U}^{\ast}}}c(o_{u^{\prime}i^{\ast}})$ by adjusting $\bm{o}_{\mathcal{N}_{\mathcal{U}^{\ast}}i^{\ast}}$. Here $c(o_{u^{\prime}i^{\ast}})$ is a cost function measuring the damage of $o_{u^{\prime}i^{\ast}}$ to neighbor $u^{\prime}$ through the decrease in $u^{\prime}$’s expected rating, with further details provided in Section 4.3.
4 METHOD
In this section, we introduce NIRec as shown in Figure 3, which consists of an interference representation-based potential outcome estimation module and a post-learning-based treatment value optimization module to achieve PRSN.
Figure 3: NIRec. Initially, train an interference representation-based estimation module for potential outcome estimation. Then, optimize the exposure of the target item to the user’s neighbors, given the well-trained estimation module and a predesigned cost function.
4.1 Identification
We identify interference-aware potential outcomes with standard assumptions in the causal inference literature [Ogburn2014caudiag, ma2021interference, ma2022hypergraph].
Assumption 1
There exists an interference representation function $\phi(\cdot)$ sufficiently capturing information of $(u,i)$’s neighbors’ treatment values $\bm{o}_{\mathcal{N}_{u}i}$ and features $\bm{\mathrm{X}}_{\mathcal{N}_{u}i}$. If $\phi(\cdot)$ is determined, $(u,i)$’s potential outcome is determined. Formally, for any $(u,i)$, $\bm{o}^{1}_{\mathcal{N}_{u}i}$, $\bm{o}^{2}_{\mathcal{N}_{u}i}$, $\bm{\mathrm{X}}^{1}_{\mathcal{N}_{u}i}$, and $\bm{\mathrm{X}}^{2}_{\mathcal{N}_{u}i}$, if $\phi(\bm{o}^{1}_{\mathcal{N}_{u}i},\bm{\mathrm{X}}^{1}_{\mathcal{N}_{u}i})=\phi(\bm{o}^{2}_{\mathcal{N}_{u}i},\bm{\mathrm{X}}^{2}_{\mathcal{N}_{u}i})$, we have:
$Y_{ui}\left(O_{ui}=1,\bm{O}_{\mathcal{N}_{u}i}=\bm{o}^{1}_{\mathcal{N}_{u}i}\right)=Y_{ui}\left(O_{ui}=1,\bm{O}_{\mathcal{N}_{u}i}=\bm{o}^{2}_{\mathcal{N}_{u}i}\right).$ (2)